two identical metal sphere carry charges +q and -2q .when the spheres are separated by a large distance r.the force btw them is F .now the sphere are allowed to touch and then moved back to the same separation .find new F
1
Expert's answer
2017-11-24T14:35:06-0500
The force between +q and -2q at a distance r is given as F
Now F= K ((+q)×(-2q))/r^2 = -2kq^2/r^2
When the chages +q and –2q comes in contact, than it neutralises the chage +q by –q charge and lefts only –q. These -q charge will be equally distributed both the sphare. If it is seperated the two sphare again than each sphare will aquire a mount of charge of –q/2 and –q/2 .
Now the new force between them will be

F_n= K ((-q^2/2)(-q^2/2))/r^2 = 1/4((kq^2)⁄r^2 ) = (-F)/8
So the force will be 1/8 part of previous force and it will be repulsive foce while F was attractive.